# Voting-Based Semi-Parallel Proof-of-Work Protocol

### v3

@@ IV-A2 Confirmation Rule
IV-A3 Numerical Evaluation
To illustrate the strength of our protocol in Bitcoin regime, we pick the following parameters for numerical evaluation of our theoretical results. We pick $\lambda_{B}=\frac{1}{600}$, $\Delta=10$ seconds [43] as usual for Bitcoin and $\alpha=0.25$ fraction of adversarial power. We further pick $\delta=1$ second. Notice that, $\delta=1$ second upper bound is on the delay of the transmission of a proof that has less than $1KB$ size is extremely pessimistic. Even though we assumed $L\geq 10$, to display the behavior of the curve fully we plot the double-spending curve for $L\geq 1$ in Figure 4. It is easy to verify that our protocol achieves an upper bound of double-spending probability around $10^{-4}$ before $L=50$.
@@ IV-B Incentive Attacks
We believe, it is trivial to do such an analysis, as it is done via Monte Carlo simulations in [4, Section V, Fig. 7], which essentially shows that PoW protocols become more resistant to entire block withholding as the number of proofs required for a single block increases. Hence, we omit such a trivial confirmation of expected results here.
IV-B1 Proof Withholding
The latter attack is what we call proof withholding attack in this paper and it is generally much harder to prevent in parallel PoW protocols. The analysis and discussion we provided in Section II shows that the existing parallel PoW protocols fail to prevent such attacks for even small values of $\alpha$ and they perform worse than Nakamoto consensus in general. Our protocol on the other hand, is a hybrid protocol where a block consists of multiple proofs like parallel PoW [4, 29, 30] and multistage PoW protocols [23, 26, 39], however the order of these proofs are semi-sequential, i.e., we allow at most two parallel proofs at a proof height which is inspired by the $k$-cluster rule in PHANTOM protocol [21] with $k=1$.
@@ IV-B1 Proof Withholding
IV-B2 Numerical Evaluation
In Appendix F, we explain the MDP model with the appropriate state and action space as well as the objective function for our protocol. In Figure 5, we present the results of the MDP models together with the results of Section II where we analyzed the relative rewards of the existing parallel PoW protocols. We pick $\gamma=0.5$ for the MDP models and truncate the results of Section II before they rise above the honest behavior for the sake of neatness in the figures. For Nakamoto consensus, we run the MDP model of [5] and the modified model described in in Appendix F is run for our protocol with $10^{-5}$ precision where $\alpha$ values range from $0.02$ to $0.48$ with a $0.02$ spacing between them.
Note that the effect of $\gamma$ is ignored in the analysis of Section II for Tailstorm and DAG-style voting, which implies that the actual relative rewards for the adversary is underestimated and the best possible rewards are even higher than the values we show here. Nevertheless, it is clear that our protocol performs better even when the adversary is underestimated in other protocols.
In our protocol, matching always gives a half reward when $(a,h,fork)=(1,1,relevant)$ (see Appendix F) whereas the adversarial reward in Nakamoto consensus increases with increasing $\gamma$ since matching succeeds more often. Thus, in comparison to the Nakamoto consensus, our protocol is more resilient against selfish mining attacks when $\gamma\geq 0.3$ and resilience increases even more as $\gamma$ increases further, which we plot in Appendix F. On the other hand, when $\alpha>0.29$ as $\gamma$ drops below $0.3$, Nakamoto consensus becomes slightly more resilient. For the relative ratio in other values of $\gamma$, we refer the reader to Appendix F.
@@ IV-B3 Reward/Punish
IV-C Transaction Fee Split Rule
@@
V Relation to existing works and limitations
@@
Note that, completely ignoring parallel proofs or accepting but not rewarding them would be against the original intention of creating a parallel PoW protocol that is fair and has less variance and allows proofs to be mined in parallel. On the other hand, allowing too many parallel proofs allows the adversary to mount withholding attacks as considered in Section II. Moreover, it is also unnecessary to allow too many parallel proofs if proof sizes are very small that can be propagated through the network fairly quick as in our protocol. Thus, our pick of $k=1$ is a compromise between them. Future work could consider fair rewarding mechanisms while allowing higher $k$-cluster proof DAGs, where $k$ can be seen as a measure of parallelness.
@@ A-A DoS Attack in Bobtail
A-B Proof Withholding in Bobtail
The authors of Bobtail also consider Monte Carlo simulations of proof withholding attacks where a miner simply withholds all its proofs for the current block height until it finds the lowest hash valued proof or gives up the attack and releases all its proofs if it seems unlikely to do so before honest miners creating a block with all proofs honest [4, Section VIII.B]. However, apparently, the authors do not consider that if the adversary manages to mine the proof with the lowest hash value for the current block height, it can release this lowest hash proof while withholding the rest of the proofs, to build a lead for the proofs of the next blocks. This in turn becomes similar to DoS attack described earlier, since honest miners will have to find a proof with lower hash value. We describe this hard-coded attack next and provide simulations to display the effectiveness of the attack.
A-C DoS Proof Withholding Attack
@@
  1. 1.
  2. 2.
    When $L$ public proofs are completed (L1L-1 honest and $1$ adversarial):
    • If the honest miners already produced a proof with lower than the public adversarial proof, they can create a block. To prevent that, the adversary releases all its work for that block plus some additional work from the lead that is has created for the next block to undo all honest work. It simply loses no reward and undoes some honest work.
@@
Notice that, to succeed further in this attack, the adversary can further modify the attack to keep its proofs hidden (Step 2) only if the smallest hash it has produced is below a certain threshold. It could simply minimize its risk by picking a low enough smallest hash region mentioned in [20]. We refer the reader to [20] to see why SHTB rule is a weak choice in consensus models against incentive attacks.
For example, in Figure 1, if $p_{5}$ is mined by the adversary and the adversary is confident that the hash value of $p_{5}$ is low enough (in smallest hash region), it can refuse to share the block it creates and it can safely mine for the next block before honest miners can do so for a while, gaining an advantage. Although the authors claim that DoS attacks can be thwarted by banning the attackers, such a defense mechanism is not straightforward in permissionless blockchain settings with easy-to-obtain sybil identities.
@@
Let $H_{waste}$ denote the number of honest proofs that has to be created until honest miners hit a lower hash value. Now, assuming the adversary has created the lowest hash value $x$,
@@
Now considering the distribution of $x$, we have
@@
which diverges. This essentially means that the adversary is able to hit so low hash values that the honest miners have to waste virtually infinite amount of energy to replace it. From a practical perspective, if the block is not created after the first $L$ public proofs, honest miners could start over with ignoring the adversarial lowest hash proof. In this case, they will start to mine until they create $L$ new proofs or a proof with lower hash than the adversarial, whichever comes first. Thus, $H_{waste}$ becomes essentially the expectation of a truncated geometric distribution capped at $(L+N)$, where $N$ denotes the number of hidden adversarial proofs that were created for the current block before the adversary started to build a lead for the next block. For the first block the adversary gets the lowest hash proof, we have,
$\mathbb{E}$ $[H_{waste}]=\int_{0}^{1}L(1-x)^{L-1}\times$
$\left(\sum_{k=1}^{L+N-1}k(1-x)^{k-1}x+(L+N)(1-x)^{L+N-1}\right)\,dx$ (6)
@@
Results with the practical rule is presented in Figure 6, where we simulate the attack for $1000000$ blocks for $L=10,25,50$. Even with the extra protection that we add with this rule, it is clear that the profitability threshold is zero and the adversarial relative reward is at par with the upper bound of selfish mining attack in Nakamoto consensus, i.e., when $\gamma=1$. As $\alpha$ rises above $\frac{1}{3}$, the attack becomes even more powerful than the upper bound of selfish mining attack in Nakamoto consensus. Moreover, without the practical rule we introduced, our simulations do verify our theory that the explosion happens when we simulate a sufficiently large number of blocks.
@@ Appendix B Why Relative Reward Metric?
More specifically, an incentive compatible system that is safe against the attack should have
@@
On the other hand, if the system is secure under the attacks (as [29, 30] claim) and honest miners are not leaving the system in an open economy model, it implies that the coin exchange value compared to their mining costs dynamically changed and honest mining became sustainable, which in turn, implies that the adversary is making unfair profits as it gets more than its fair share. In other words, under the attack, if the honest reward per chain progress per computational power, $\frac{\nu_{attack}}{1-\alpha}$ is sustainable when taking the mining costs into the account, then adversarial reward per chain progress per computational power is sustainable if $\frac{\mu_{attack}}{\alpha}=\frac{\nu_{attack}}{1-\alpha}$. Since $\rho>\alpha$, the adversarial reward per chain progress per computational power is not only sustainable but also more profitable, i.e.,
@@
What [29, 30] do, is decreasing the value of $c$ when the system is under the attack and focusing on the loss of the adversary $\mu_{attack}$, completely ignoring a potentially larger and disastrous loss on the honest side $\nu_{attack}$ and claiming their system is secure. Since [29, 30] assume that the system is secure under the attacks, which implies that honest miners are rewarded enough to continue on mining, and then they should focus on the adversarial rewards to see if they get more than their fair share, i.e., check if (9) is satisfied or not, which is equivalent to checking if $\rho\leq\alpha$ or not.
@@ Appendix C Proof Withholding Attack for Tailstorm and DAG-style Voting
Figure 8: DAG version of the withholding attack
@@
With the hard-coded proof withholding attack, the fraction of the adversarial proofs in the public chain reaches $\frac{\alpha}{1-\alpha}\frac{L-1}{L}$ even when $\gamma=0$, for both tree- and DAG-style voting. Note that, in Tailstorm [29], although the reward a block gets is scaled by the depth of the proof tree, everyone is punished and rewarded equally. Hence, the fraction of adversarial proofs we derive, is the fraction of the rewards the adversary gets. We note that the authors of [29] consider a protocol variant, TS/const, where the relative reward of the adversary equals reward per chain progress. Their simulations include an incentive attack called minor delay against TS/const, which seems similar to the attack we describe here. However, the claims of the authors in [29, Figure 4] regarding the attack are not verifiable given our theoretical analysis. Hence, in Appendix C-B, we provide simulations of a practical implementation of the attack for $100000$ blocks and verify our theoretical results.
C-A DAG Version of the Attack
@@
$\mathbb{E}[A_{R}]=$ $1\cdot L+(L\mathbb{E}[b_{n}]-1)\cdot L\mathbb{E}[a_{n}]$
$+L(\mathbb{E}[a_{n}]-\mathbb{E}[b_{n}])\cdot L\frac{1+\mathbb{E}[a_{n}]}{2},$ (17)
@@ C-B Simulations
The results are provided in Figure 9, where we run the simulation for $100000$ blocks for each parameter. As we mentioned in our theoretical analysis, the accuracy of our analysis depends on $L$ as we use expected values. Hence, as $L$ increases, the theoretical and simulation values converge. Further, the attack does not need any network advantage $\gamma$ and both protocols become less resistant to incentive attacks than the original Nakamoto consensus. We note that, we did not dig through all possible ways to find better practical implementations for which the simulation results could converge faster even for smaller $L$. We leave this to interested readers, since our purpose is to show the vulnerability of the existing parallel PoW protocols.
@@ Appendix D Fork Choice Rule
On the other hand, if an honest miner owns the ledger offering the most fee and the adversary creates the first initiator proof for the next block height, honest miners will accept the adversarial initiator proof since it is the chain with most aggregate work. Notice, the adversarial initiator proof might refer to another valid ledger instead of the ledger offering the most fee. Such a situation does not create any issue from the perspective of the consensus since the the fork choice rule is always the same, i.e., pick the chain that contains the most aggregate work. From the perspective of the transaction fee revenues, in Section IV-C and Appendix G we show that such an attack is not profitable for the adversary.
@@ E-1 The Attack
We first remind the reader that a double-spending analysis is more complicated than simply analyzing a private attack when network delays are present. In other words, when network delays are present, the view of honest miners differ and some honest miners can be lured into working for the adversary temporarily during delays, increasing its hashrate, known as balanced attacks [40, Lemma F. 1]. This in turn complicates the analysis.
Without network delays, a private attack can be shown to succeed if any other strategy succeeds in Nakamoto consensus [40, Theorem 5.1]. This result is easily extendable to the heaviest chain rule here if we had no network delays. To analyze double-spending under network delays, we will give the adversary certain advantages that simplify our analysis and give an upper bound for the double-spending probability. In other words, we will provide achievable consistency results from the honest miners perspective.
@@
E-2 Analysis
Assume at time $\tau$, the block $h$ has $L$ proofs and a ledger of these proofs are starting to be disseminated through network. At this point, the honest miners need $\Delta$ time to receive an honest ledger of block $h$. From this point onward, honest miners will have to create a complete block at height $h+1$ with $L$ (or $L+1$) proofs each of which will experience a $\delta$ delay. When these proofs are mined and dispersed, another $\Delta$ time is needed for the proposed ledger of block $h+1$ to be received by honest miners. We call the time the ledger of block $h+1$ is received by all honest miners as $\tau_{c}$. Notice that, at this point, the ledger of the block $h$ is confirmed according to $1$-block confirmation rule since block $h+1$ is complete. Thus, the interval $\tau,\tau_{c}]$ is called confirmation interval.

### v4

@@ 4.1.2 Confirmation Rule
4.1.3 Numerical Evaluation
To illustrate the strength of our protocol in the Bitcoin regime, we pick the following parameters for numerical evaluation of our theoretical results. We pick $\lambda_{B}=\frac{1}{600}$, $\Delta=10$ seconds [43] as usual for Bitcoin and $\alpha=0.25$ fraction of adversarial power. We further pick $\delta=1$ second. Notice that, $\delta=1$ second upper bound is on the delay of the transmission of a proof that has less than $1KB$ size is extremely pessimistic. Even though we assumed $L\geq 10$, to display the behavior of the curve fully we plot the double-spending curve for $L\geq 1$ in Figure 4. It is easy to verify that our protocol achieves an upper bound of double-spending probability around $10^{-4}$ before $L=50$.
@@ 4.2 Incentive Attacks
We believe, it is trivial to do such an analysis, as it is done via Monte Carlo simulations in [4, Section V, Fig. 7], which essentially shows that PoW protocols become more resistant to entire block withholding as the number of proofs required for a single block increases; hence, we omit such trivial confirmation of expected results here.
4.2.1 Proof Withholding
The latter attack is what we call proof withholding attack in this paper and it is generally much harder to prevent in parallel PoW protocols. The analysis and discussion we provided in Section 2 show that the existing parallel PoW protocols fail to prevent such attacks for even small values of $\alpha$ and they perform worse than Nakamoto consensus in general. Our protocol on the other hand, is a hybrid protocol where a block consists of multiple proofs like parallel PoW [4, 29, 30] and multistage PoW protocols [23, 26, 39], however the order of these proofs are semi-sequential, i.e., we allow at most two parallel proofs at a proof height which is inspired by the $k$-cluster rule in PHANTOM protocol [21] with $k=1$.
@@ 4.2.1 Proof Withholding
4.2.2 Numerical Evaluation
In Appendix F, we explain the MDP model with the appropriate state and action space as well as the objective function for our protocol. In Figure 5, we present the results of the MDP models together with the results of Section 2 where we analyzed the relative rewards of the existing parallel PoW protocols. We pick $\gamma=0.5$ for the MDP models and truncate the results of Section 2 before they rise above the honest behavior for the sake of neatness in the figures. For Nakamoto consensus, we run the MDP model of [5] and the modified model described in in Appendix F is run for our protocol with $10^{-5}$ precision where $\alpha$ values range from $0.02$ to $0.48$ with a $0.02$ spacing between them.
Note that the effect of $\gamma$ is ignored in the analysis of Section 2 for Tailstorm and DAG-style voting, which implies that the actual relative rewards for the adversary is underestimated and the best possible rewards are even higher than the values we show here. Nevertheless, it is clear that our protocol performs better even when the adversary is underestimated in other protocols.
In our protocol, matching always gives a half reward when $(a,h,fork)=(1,1,relevant)$ (see Appendix F) whereas the adversarial reward in Nakamoto consensus increases with increasing $\gamma$ since matching succeeds more often. Thus, in comparison to the Nakamoto consensus, our protocol is more resilient against selfish mining attacks when $\gamma\geq 0.3$ and resilience increases even more as $\gamma$ increases further, which we plot in Appendix F. On the other hand, when $\alpha>0.28$ as $\gamma$ drops below $0.3$, Nakamoto consensus becomes slightly more resilient. For the relative ratio in other values of $\gamma$, we refer the reader to Appendix F.
@@ 4.2.3 Reward/Punish
4.3 Transaction Fee Split Rule
@@
5 Relation to Existing Works and Limitations
@@
Note that, completely ignoring parallel proofs or accepting but not rewarding them would be against the original intention of creating a parallel PoW protocol that is fair and has less variance and allows proofs to be mined in parallel. On the other hand, allowing too many parallel proofs allows the adversary to mount withholding attacks as considered in Section 2. Moreover, it is also unnecessary to allow too many parallel proofs if proof sizes are very small that can be propagated through the network fairly quick as in our protocol. Thus, our pick of $k=1$ is a compromise between them. Future work could consider fair rewarding mechanisms while allowing higher $k$-cluster proof DAGs, where $k$ can be seen as a measure of parallelness.
@@ A.1 DoS Attack in Bobtail
A.2 Proof Withholding in Bobtail
The authors of Bobtail also consider Monte Carlo simulations of proof withholding attacks where a miner simply withholds all its proofs for the current block height until it finds the lowest hash valued proof or gives up the attack and releases all its proofs if it seems unlikely to do so before honest miners creating a block with all proofs honest [4, Section VIII.B]. However, the authors do not consider that if the adversary manages to mine the proof with the lowest hash value for the current block height, it can release this lowest hash proof while withholding the rest of the proofs, to build a lead for the proofs of the next blocks. This in turn becomes similar to the DoS attack described earlier, since honest miners will have to find a proof with lower hash value. We describe this hard-coded attack next and provide simulations to display the effectiveness of the attack.
A.3 DoS Proof Withholding Attack
  1. 1.
  2. 2.
    When $L$ public proofs are completed (L1L-1 honest and $1$ adversarial):
    • If the honest miners already produced a proof with lower than the public adversarial proof, they can create a block. To prevent that, the adversary releases all its work for that block plus some additional work from the lead that it has created for the next block to undo all honest work. It simply loses no reward and undoes some honest work.
@@ A.3 DoS Proof Withholding Attack
Notice that, to succeed further in this attack, the adversary can further modify the attack to keep its proofs hidden (Step 2) only if the smallest hash it has produced is below a certain threshold. It could simply minimize its risk by picking a low enough smallest hash region mentioned in [20]; see [20] to see why SHTB rule is a weak choice in consensus models against incentive attacks.
For example, in Figure 1, if $p_{5}$ is mined by the adversary and the adversary is confident that the hash value of $p_{5}$ is low enough (in smallest hash region), it can refuse to share the block it creates and it can safely mine for the next block before honest miners can do so for a while, gaining an advantage. Although the authors state that DoS attacks can be thwarted by banning the attackers, such a defense mechanism is not straightforward in permissionless blockchain settings with easy-to-obtain sybil identities.
@@
Let $H_{waste}$ denote the number of honest proofs that has to be created until the honest miners hit a lower hash value. Now, assuming the adversary has created the lowest hash value $x$,
@@
Now, considering the distribution of $x$, we have
@@
which diverges. This essentially means that the adversary is able to hit so low hash values that the honest miners have to waste virtually infinite amount of energy to replace it. From a practical perspective, if the block is not created after the first $L$ public proofs, honest miners could start over with ignoring the adversarial lowest hash proof. In this case, they will start to mine until they create $L$ new proofs or a proof with lower hash than the adversarial, whichever comes first. Thus, $H_{waste}$ becomes essentially the expectation of a truncated geometric distribution capped at $(L+N)$, where $N$ denotes the number of hidden adversarial proofs that were created for the current block before the adversary started to build a lead for the next block. For the first block the adversary gets the lowest hash proof, we have
$\mathbb{E}[H_{waste}]$ $=\int_{0}^{1}L(1-x)^{L-1}\times\left(\sum_{k=1}^{L+N-1}k(1-x)^{k-1}x+(L+N)(1-x)^{L+N-1}\right)\,dx$ (6)
@@
Results with the practical rule is presented in Figure 6, where we simulate the attack for $10^{6}$ blocks for $L=10,25,50$. Even with the extra protection that we add with this rule, it is clear that the profitability threshold is zero and the adversarial relative reward is at par with the upper bound of selfish mining attack in Nakamoto consensus, i.e., when $\gamma=1$. As $\alpha$ rises above $\frac{1}{3}$, the attack becomes even more powerful than the upper bound of selfish mining attack in Nakamoto consensus. Moreover, without the practical rule we introduced, our simulations do verify our theory that the explosion happens when we simulate a sufficiently large number of blocks.
@@ Appendix B Why Relative Reward Metric?
Specifically, an incentive compatible system that is safe against the attack should have
@@
On the other hand, if the system is secure under the attacks (as [29, 30] claim) and honest miners are not leaving the system in an open economy model, it implies that the coin exchange value compared to their mining costs dynamically changed and honest mining became sustainable, which in turn, implies that the adversary is making unfair profits as it gets more than its fair share. In other words, under the attack, if the honest reward per chain progress per computational power, $\frac{\nu_{attack}}{1-\alpha}$ is sustainable when taking the mining costs into account, then adversarial reward per chain progress per computational power is sustainable if $\frac{\mu_{attack}}{\alpha}=\frac{\nu_{attack}}{1-\alpha}$. Since $\rho>\alpha$, the adversarial reward per chain progress per computational power is not only sustainable but also more profitable, i.e.,
@@
What [29, 30] do, is decreasing the value of $c$ when the system is under the attack and focusing on the loss of the adversary $\mu_{attack}$, ignoring a potentially larger and disastrous loss on the honest side $\nu_{attack}$ and claiming their system is secure. Since [29, 30] assume that the system is secure under the attacks, which implies that honest miners are rewarded enough to continue on mining, and then they should focus on the adversarial rewards to see if they get more than their fair share, i.e., check if (9) is satisfied or not, which is equivalent to checking if $\rho\leq\alpha$ or not.
@@ Appendix C Proof Withholding Attack for Tailstorm and DAG-Style Voting
Figure 8: DAG version of the withholding attack.
@@
With the hard-coded proof withholding attack, the fraction of the adversarial proofs in the public chain reaches $\frac{\alpha}{1-\alpha}\frac{L-1}{L}$ even when $\gamma=0$, for both tree- and DAG-style voting. Note that, in Tailstorm [29], although the reward a block gets is scaled by the depth of the proof tree, everyone is punished and rewarded equally. Hence, the fraction of adversarial proofs we derive, is the fraction of the rewards the adversary gets. We note that the authors of [29] consider a protocol variant, TS/const, where the relative reward of the adversary equals reward per chain progress. Their simulations include an incentive attack called minor delay against TS/const, which seems similar to the attack we describe here. However, the claims of the authors in [29, Figure 4] regarding the attack are not verifiable given our theoretical analysis. Hence, in Appendix C.2, we provide simulations of a practical implementation of the attack for $10^{5}$ blocks and verify our theoretical results.
C.1 DAG Version of the Attack
@@
$\mathbb{E}[A_{R}]=$ $1\cdot L+(L\mathbb{E}[b_{n}]-1)\cdot L\mathbb{E}[a_{n}]+L(\mathbb{E}[a_{n}]-\mathbb{E}[b_{n}])\cdot L\frac{1+\mathbb{E}[a_{n}]}{2},$ (17)
@@ C.2 Simulations
The results are provided in Figure 9, where we run the simulation for $10^{5}$ blocks for each parameter. As we mentioned in our theoretical analysis, the accuracy of our analysis depends on $L$ as we use expected values. Hence, as $L$ increases, the theoretical and simulation values converge. Further, the attack does not need any network advantage $\gamma$ and both protocols become less resistant to incentive attacks than the original Nakamoto consensus. We note that, we did not dig through all possible ways to find better practical implementations for which the simulation results could converge faster even for smaller $L$. We leave this to interested readers, since our purpose is to show the vulnerability of the existing parallel PoW protocols.
@@ Appendix D Fork Choice Rule
On the other hand, if an honest miner owns the ledger offering the most fee and the adversary creates the first initiator proof for the next block height, honest miners will accept the adversarial initiator proof since it is the chain with most aggregate work. Notice, the adversarial initiator proof might refer to another valid ledger instead of the ledger offering the most fee. Such a situation does not create any issue from the perspective of the consensus since the fork choice rule is always the same, i.e., pick the chain that contains the most aggregate work. From the perspective of the transaction fee revenues, in Section 4.3 and Appendix G, we show that such an attack is not profitable for the adversary.
@@ E.1 The Attack
We first recall that a double-spending analysis is more complicated than simply analyzing a private attack when network delays are present. In other words, when network delays are present, the views of honest miners differ and some honest miners can be lured into working for the adversary temporarily during delays, increasing its hashrate, known as balanced attacks [40, Lemma F. 1]. This in turn complicates the analysis.
Without network delays, a private attack can be shown to succeed if any other strategy succeeds in Nakamoto consensus [40, Theorem 5.1]. This result is easily extendable to the heaviest chain rule here if we had no network delays. To analyze double-spending under network delays, we will give the adversary certain advantages that simplify our analysis and give an upper bound for the double-spending probability. In other words, we will provide achievable consistency results from the honest miners’ perspective.
@@
E.2 Analysis
Assume at time $\tau$, the block $h$ has $L$ proofs and a ledger of these proofs are starting to be disseminated through the network. At this point, the honest miners need $\Delta$ time to receive an honest ledger of block $h$. From this point onward, honest miners will have to create a complete block at height $h+1$ with $L$ (or $L+1$) proofs each of which will experience a $\delta$ delay. When these proofs are mined and dispersed, another $\Delta$ time is needed for the proposed ledger of block $h+1$ to be received by honest miners. We call the time the ledger of block $h+1$ is received by all honest miners as $\tau_{c}$. Notice that, at this point, the ledger of the block $h$ is confirmed according to $1$-block confirmation rule since block $h+1$ is complete. Thus, the interval $\tau,\tau_{c}]$ is called confirmation interval.